### v2

@@ Methods
Table 3: Enzyme-class guidance results for different EC sets at 80% masking level. Number of samples (N) per set are indicated. “Mask Level %” is the percent of positions masked, 100 % is full masked. “CLEAN %” is the percentage of samples that were correctly classified as the intended target class. “pLDDT $>0.8$” is the proportion of samples correctly reclassified by CLEAN that have a median pLDDT across residues above 0.8.
Method EC Set CLEAN % ($\uparrow$) Mean pLDDT ($\uparrow$) pLDDT $>0.8$ ($\uparrow$) Success % ($\uparrow$) N
@@
Table 4: Test set accuracies of the fold classifier used for evaluation in CATH-guided structure generation. Accuracies are computed for each of the CATH level separately. Both Micro and Macro accuracies are reported as alternative ways to average over the different classes within each CATH level [74].
Level Micro Accuracy ($\uparrow$) Macro Accuracy ($\uparrow$)
C 0.9888 0.9359
A 0.9678 0.9702
T 0.9490 0.9212
Table 5: CATH classes used for conditioning in the fold class guidance experiment, grouped by CATH level and frequency of occurrence in the AFDB. The frequency groups are defined following Geffner et al. [18]. For T level classes, 10 classes are randomly sampled for each of the frequency groups.
CATH Level Frequency Classes
A Level Common (n=9) 2.40, 3.40, 1.20, 1.10, 3.90, 3.30, 3.10, 3.20, 2.60
Regular (n=13) 2.30, 1.25, 2.170, 4.10, 3.60, 2.20, 3.55, 3.50, 2.130, 2.10, 2.160, 2.70, 2.102
Rare (n=17) 2.120, 3.15, 2.50, 2.80, 2.100, 1.50, 3.80, 1.40, 3.65, 2.115, 3.75, 3.100, 3.70, 2.110, 2.150, 2.140, 2.90
T Level Common (n=10) 3.40.190, 2.30.110, 3.40.50, 3.30.70, 3.10.20, 1.10.287, 2.60.40, 1.25.40, 3.10.129, 2.30.30
Regular (n=10) 1.10.1510, 3.30.1060, 3.90.1030, 1.20.1440, 2.160.10, 3.40.225, 3.30.530, 3.30.429, 3.60.15, 3.10.310
Rare (n=10) 3.100.10, 2.40.155, 3.30.2430, 3.40.605, 3.30.2370, 1.20.91, 3.90.1470, 3.20.130, 3.90.290, 3.20.190
Table 6: Illumina PCR1 primers
@@
Supplementary figures
Figure 3: Class balanced accuracy of the enzyme classifier on test sequences from the 100 most common level-4 EC classes at different masking levels. The x-axis shows the masking level from fully masked (left) to fully unmasked (right), while the y-axis shows the balanced accuracy for each of the four EC hierarchical levels.
Figure 4: Mean pLDDT scores across all samples in the top-10 EC classes experiment. The plot compares mean pLDDT scores between guided and unconditional generation methods across different masking levels (50%-100%). Higher pLDDT scores indicate better structural confidence in the generated protein sequences, with guided generation consistently producing higher quality structures.
Figure 5: Additional fold class generation results with varying sampling temperatures for either ESM3 or ESM3 with ProteinGuide. The x-axis shows the sampling temperature used with ESM3 when generating backbone structures. The height of each bar shows the percentage of structures correctly classified to their target fold class, while the hatched sub-portion indicates the subset that are also designable (self-consistency RMSD $\leq 2$Å). Results are grouped by CATH level (row). For level C (top row), there are three classes: mainly $\alpha$ (left), mainly $\beta$ (middle), and mixed $\alpha/\beta$ (right). For level A (middle row) and T (bottom row), we grouped the classes for each level by their frequency of occurrence in the AFDB into Common (left), Regular (middle), and Rare (right). The metrics were aggregated within each group.
Figure 6: Diversity of generated backbone structures for C level, with either ESM3 or ESM3 with ProteinGuide. Diversity is quantified with TM-scores, which range from zero to one, where lower scores indicate higher diversity. For each C level classes (mainly $\alpha$, mainly $\beta$, and mixed $\alpha/\beta$), 10,000 structures were sampled, either using unguided ESM3 (teal) or with ProteinGuide(pink). The generated samples were filtered to those that are both classified correctly to the desired fold class by the fold classifier and are designable (scRMSD $\leq 2$Å). These samples were further stratified by length, and the pairwise TM-score was computed among all sampled structures for each combination of class and length. The markers on each line indicate the averaged pairwise TM-score for specific combination of class and length. Error bands represent standard deviation of pairwise TM scores within each length group. The annotated numbers indicate the number of structures that are both classified correctly and designable for each class and length.
Figure 7: Examples of generated protein backbone structures by guiding ESM3 with ProteinGuide to condition on selected examples CATH classes. The shown proteins are both designable and correctly classified into the desired conditioning fold class. Numbers in parentheses are the corresponding CATH code for the labeled fold classes. Two level digits (e.g., 2.40) represent A level conditioning, and three level digits (e.g., 3.20.20) represent T level conditioning.
@@
Figure 8: Experimental characterization of sequences sampled without ProteinGuide. a. Log enrichment scores of the $2,000$ sequences sampled from ESM3 and FMIF for two biological replicates. Each sequence is colored by the number of mutations away from TadA (red). The low vs. high activity threshold used to train a classifier for guidance is shown as a dotted grey line. b. The same log enrichment scores as in (a.) but different colors and markers are used to denote which model the sequences were generated from.
@@
Figure 9: Experimental characterization of sequences sampled with ProteinGuide. a. Log enrichment scores of the $1,888$ sequences sampled from ESM3 and FMIF with ProteinGuide for two biological replicates. Each sequence is colored by the number of mutations away from TadA (red). b. The same log enrichment scores as in (a.) but different colors and markers are used to denote which model the sequences were generated from.
@@
Figure 10: Normalizing log enrichment scores between round 1 and round 2 a. Log enrichment scores of the $112$ control sequences that were tested in both round 1 and round 2. The x-axis shows the median log enrichment scores for the replicates from round 1. The y-axis shows the median log enrichment scores for the replicates from round 2. b. Histogram of log enrichment values comparing unconditional sequences (teal) and guided sequences (pink) after calibration between rounds. Bootstrap analysis (1,000 iterations) of the 112 control sequences yielded a mean median log enrichment of 0.56 for the guided library with a 90% confidence interval of [0.38, 0.75]. The 5th and 95th percentile mappings (gold and gray histograms, respectively) represent conservative and optimistic transformations based on bootstrap sampling, visualizing the uncertainty range in our calibration procedure. The gray dashed line at a log enrichment value of 0 represents the low vs. high activity classification threshold. Our guidance strategy was designed to generate sequences with log enrichment values greater than this threshold.
@@
Figure 11: Base editor (starting variant) protein sequence
Figure 12: Reporter gene DNA sequence

### v4

@@ 3 Supplementary Tables
Table S5: Enzyme-class guidance results for different EC sets at 80% masking level. Number of samples (N) per set are indicated. “Mask Level %” is the percent of positions masked, 100 % is full masked. “CLEAN %” is the percentage of samples that were correctly classified as the intended target class. “pLDDT >0.8>0.8” is the proportion of samples correctly reclassified by CLEAN that have a median pLDDT across residues above 0.8.
Method EC Set CLEAN % ($\uparrow$) Mean pLDDT ($\uparrow$) pLDDT >0.8>0.8 ($\uparrow$) Success % ($\uparrow$) N
@@
Table S6: CATH labels used for A- and T-level fold-class conditioning baselines. Each row includes the CATH code, its description, whether it belongs to the “common” or “random” group, the total number of samples in the dataset, and the fold-oracle test accuracy for that label.
Code Description Group # samples Test Acc
A-level
3.40 3-Layer(aba) Sandwich common 84172 0.8163
2.60 Sandwich common 34728 0.8724
2.40 Beta Barrel common 24208 0.7961
3.20 Alpha-Beta Barrel common 15623 0.7755
3.10 Roll common 13951 0.7230
2.30 Roll common 6424 0.7406
1.25 Alpha Horseshoe common 2967 0.7316
2.80 Trefoil common 1365 0.8224
2.130 7 Propeller common 1239 0.9423
2.120 6 Propeller common 1011 0.7374
2.100 Aligned Prism random 305 0.8485
3.80 Alpha-Beta Horseshoe random 756 0.8226
2.102 3-layer Sandwich random 234 0.7812
2.50 Clam random 66 0.9091
3.70 Box random 297 0.9412
2.115 5 Propeller random 434 0.7778
2.90 Orthogonal Prism random 141 1.0000
2.160 3 Solenoid random 964 0.7396
3.65 Alpha-beta prism random 400 1.0000
2.140 8 Propeller random 398 0.8889
T-level
3.40.50 Rossmann fold common 50985 0.7885
2.60.40 Immunoglobulin-like common 23273 0.8811
3.20.20 TIM Barrel common 14701 0.7779
3.30.70 Alpha-Beta Plaits common 9911 0.7960
2.60.120 Jelly Rolls common 8713 0.8591
2.40.10 Thrombin, subunit H common 7012 0.9036
3.40.190 D-Maltodextrin-Binding Protein common 5938 0.7025
2.40.70 Cathepsin D, subunit A common 4683 1.0000
3.40.30 Glutaredoxin common 3702 0.9466
1.10.490 Globin-like common 2682 0.9792
2.30.110 Pnp Oxidase random 367 0.9487
1.25.40 Serine Threonine Protein Phosp… random 2155 0.7453
3.30.429 Macrophage Migration Inhibitor… random 489 1.0000
3.10.310 Diaminopimelate Epimerase random 285 0.8000
2.160.10 UDP N-Acetylglucosamine Acyltr… random 544 0.7547
3.10.129 Thiol Ester Dehydrase random 973 0.9024
1.20.91 Influenza Virus Matrix Protein random 76 0.7273
3.40.605 Aldehyde Dehydrogenase random 1113 1.0000
2.40.155 Green Fluorescent Protein random 1100 1.0000
3.40.225 L-fuculose-1-phosphate Aldolase random 90 1.0000
Table S7: Illumina PCR1 primers used in the base editor experiment.
@@
4 Supplementary Figures
Figure S1: Diversity and novelty of generated sequences for all methods in the ProteinMPNN stability experiment. a) Diversity of the generated sequences. For each method, diversity is evaluated on 100 sequences. Diversity is computed as the averaged pairwise hamming distance among the generated sequences. For each protein, the “Reference Set” diversity is computed on 500 randomly sampled sequences from the labeled dataset for each protein. b) Novelty of the generated sequences. For each method, novelty is evaluated on 100 sequences. For each sequence, novelty is computed as its minimum hamming distance with sequences in the labeled dataset. For each method, the mean novelty among the generated sequences is shown.
Figure S2: Datasets and oracle models used for the extrapolative experiment of guiding ESM3 with assay-informed predictive models. a) Distribution of variant counts in each dataset stratified by the hamming distance from the wild-type sequence. b) Scatterplot of the labeled vs. oracle predicted property value in each dataset, for both the training and validation set.
Figure S3: Structure prediction-based metrics for generated sequences from guiding ESM3 with assay-informed predictive models. a) Distribution of the pLDDT of the generated sequences from each method (higher is better). pLDDT is based on ESMFold prediction. Red dotted line indicates the threshold above with a generated sequence is considered a success (pLDDT $\geq 80$). b) Distribution of the refolded RMSD (Å) of the generated sequences from each method (lower is better). RMSD is computed between the predicted structure of the wild-type sequence and that of the generated sequence.
Figure S4: Multi-property, Pareto-extrapolative experiment: unfiltered generation results, heuristic constraint adjustment, and negative log-likelihood curves. a-d) Show the unfiltered compute-matched sequences generated by the three methods benchmarked in the main text and MOCE, the method used by Wang et al. [116]. e) Example of training-data-subsampled predictors (blue) against the predictor used for guidance (red). The residuals are “Predictor” minus “Subsampled”. Residuals were calculated on randomly sampled sequences of up to 30 mutations that were predicted to be in the target region for the Pareto-extrapolative task (Fig. 4b, yellow box). f) Time-series of negative log likelihoods for each method on the set of sequences predicted by the oracle to be in the target region of the Pareto-extrapolative task (Fig. 4b, yellow box). Dashed line shows the time point from which generation started. A cross entropy of 3 on the y-axis corresponds to the negative log-likelihood on the data for a uniform distribution over amino acids.
Figure S5: Data splits used in the multi-property, Pareto-extrapolative PbrR experiment. a. Summary of training and validation splits used for the PbrR experiment. Gray points show full dataset points not included in that split. The validation split was common between the predictor and finetuned model. The finetuning training data was the portion of the predictor training data that was above the dashed on-/off-target threshold line shown in the center subpanel. b. Predictions from the regression model used for ProteinGuide compared with the true assay value on the validation set.
Figure S6: Guiding ESM3 with enzyme annotations. a) Schematic for using ProteinGuide to guide ESM3 with an enzyme class predictor. b) Classification accuracy of the predictor on SwissProt test sequences as an increasing number of positions are unmasked. All accuracies are class balanced. c) Four generative methods (in legend) are assessed across 10 of the most common enzyme classes in SwissProt (horizontal axis of rightmost sub-figure). Left panel shows average success rate across all 10 common classes. Center shows the number of ECs for which each method produced at least one success. Right panel shows success rate by enzyme class for the 10 most common ECs. Top-k experiments generated 210 samples total. Out of these, the top 10 sequences rated by the predictor as most likely to be in the target EC are kept for evaluation. The height of each bar represents the percentage of sequences predicted by the oracle to be of the target EC. The hatched sub-portion are the fraction of sequences in the target EC that are also predicted to fold with pLDDT $\leq$ 0.8. Error bars are standard errors for proportions. d) Guidance performance when redesigning test set enzyme sequences. The x-axis of each subplot is the percent of positions in the original enzyme that were masked. The left subplot y-axis shows the percent of sequences generated (10 total) that were reclassified to be in the target enzyme class. The right subplot y-axis shows the percent of these 10 sequences that had a pLDDT greater than 0.8.
Figure S7: Guiding ESM3 to generate backbone structures with CATH fold class annotations. a) Schematic for using ProteinGuide to guide ESM3 with a fold class predictor (trained on CATH) to generate structures predicted to belong to specific fold class. b) Aggregated success rates of different mehods (in legends) across different CATH hierarchy levels. The height of each bar shows the percentage of structures correctly classified to their target fold class, while the hatched sub-portion indicates the subset that are also designable (self-consistency RMSD $\leq 2$Å). Results are grouped by CATH level, and the classes for each level were grouped by their frequency of occurrence in the training dataset of the fold class oracle. Metrics were aggregated within each group following Geffner et al. [35]. “Common” consists of the 10 most common classes for each level, and “Random” consists of 10 randomly selected classes for each level. c) Per-class success rates of different methods for the Common classes for both A level (top) and T level (bottom). d) Per-class success rates of different methods for the Random classes for both A level (top) and T level (bottom). e) Examples of generated backbone structures by guiding ESM3 with some example CATH classes. The shown examples are both classified correctly to the target fold class by the fold classifier and are designable.
@@ 4 Supplementary Figures
Figure S8: Experimental characterization of sequences sampled without ProteinGuide. a) Log enrichment scores of the $2,000$ sequences sampled from ESM3 and FMIF for two biological replicates. Each sequence is colored by the number of mutations away from TadA (red). The class boundary used to train a classifier is shown as a dotted grey line. b) The same log enrichment scores as in (a.) but different colors and markers are used to denote which model the sequences were generated from.
@@
Figure S9: Experimental characterization of sequences sampled with ProteinGuide. a) Log enrichment scores of the $1,888$ sequences sampled from ESM3 and FMIF with ProteinGuide for two biological replicates. Each sequence is colored by the number of mutations away from TadA (red). b) The same log enrichment scores as in a) but different colors and markers are used to denote which model the sequences were generated from.
@@
Figure S10: Normalizing log enrichment scores between round 1 and round 2. a) Log enrichment scores of the $112$ control sequences that were tested in both round 1 and round 2. The x-axis shows the median log enrichment scores for the replicates from round 1. The y-axis shows the median log enrichment scores for the replicates from round 2. b) Histogram of log enrichment values comparing unconditional sequences (teal) and guided sequences (pink) after calibration between rounds. Bootstrap analysis (1,000 iterations) of the 112 control sequences yielded a mean median log enrichment of 0.56 for the guided library with a 90% confidence interval of [0.38, 0.75]. The 5th and 95th percentile mappings (gold and gray histograms, respectively) represent conservative and optimistic transformations based on bootstrap sampling, visualizing the uncertainty range in our calibration procedure. The gray dashed line at a log enrichment value of 0 represents the low vs. high activity classification threshold. Our guidance strategy was designed to generate sequences with log enrichment values greater than this threshold.
@@
Figure S11: Supplementary sequences from the base editor experiment. a) Base editor (starting variant) protein sequence. b) Reporter gene DNA sequence.